OpenStudy (anonymous):

Evaluate: f(x) = 2x + 5, if x is 3

OpenStudy (anonymous):

Plug 3 in for x. 2(3)+5 simplify it.

OpenStudy (the_fizicx99):

f(x) = 2x + 5 f(3) = 2(3) + 5 f(3) = 6 + 5 f(3) = 11